OpenStudy (anonymous):

Simplify the expression. 2(3y - 7) - 5y(2 - y)

OpenStudy (ajprincess):

2(3y - 7) - 5y(2 - y) =2*3y-2*7-5y*2-5y*(-y) =6y-14-10y+5y^2 =5y^2-4y-14

OpenStudy (anonymous):

2(3y-7)-5y(2-y) = 6y-14 -10y +5y^2 =5y^2- 4y -14